# The Ultimate Configuration Management Tool? Lessons from a Mixed Methods Study of Ansible’s Challenges

### v2

@@ Abstract.
Infrastructure as Code (IaC) tools have transformed the way IT infrastructure is automated and managed, but their growing adoption has also exposed numerous challenges for practitioners. In this paper, we investigate these challenges through the lens of Ansible, a popular IaC tool. Using a mixed methods approach, we investigate challenges, obstacles, and issues faced by practitioners. We analyze 59,157 posts from Stack Overflow, Reddit, and the Ansible Forum to identify common pain points, complemented by 20 semi-structured interviews with practitioners of varying expertise levels.
Based on our findings, we propose four main recommendations to improve Ansible: 1) refactoring to mitigate performance issues, 2) restructuring higher-level language concepts, 3) improved debugging and error reporting tools, and 4) better documentation and learning resources. By highlighting the real-world struggles of Ansible users, we provide actionable insights for tool designers, educators, and the broader IaC community, contributing to a deeper understanding of the trade-offs inherent in IaC tools.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06copyright: none
@@ 1. Introduction
In this paper, we aim to deepen the understanding of the challenges faced by Ansible users. Ansible is currently the most widely used configuration management tool (Stack Overflow, 2024) and, although Ansible is known to have issues, such as difficulties with variable scoping (Opdebeeck, 2024) and debugging (Hacker News User gizdan, 2021), the specifics of these challenges and their impact on practitioners remain underexplored. Using a mixed methods approach, we analyze a large corpus of online discussions and conduct semi-structured interviews to systematically examine the obstacles faced by Ansible users. Our study is the first to provide a detailed account of these issues and to offer actionable recommendations for improvement.
In summary, our contributions are:
@@
  • Recommendations for improving Ansible, focusing on documentation, debugging tools, and architectural enhancements.
  • Broader implications for IaC tool the design and evolution, highlighting trade-offs between usability, flexibility, and scalability.
Our study aims to inform Ansible’s developers, educators, and the broader IaC community, fostering improvements that enhance the effectiveness and usability of IaC tools. A replication package with all the data and code used in this study is available at:
https://doi.org/10.5281/zenodo.15031088
@@ 2. Ansible
According to the Stack Overflow Developers Survey 2024, Ansible is the most widely used configuration management tool (Stack Overflow, 2024).
To define the desired system state, practitioners must write declarative code using Ansible’s hierarchical structure, which consists of four levels: roles, playbooks, plays, and tasks. Roles serve as reusable components that organize related functionality, including tasks, variables, modules, and templates. Playbooks, written in YAML, define configurations by grouping multiple plays. Each play associates specific hosts, identified by tags or hostnames, with a set of tasks. Tasks are the fundamental execution units in Ansible, consisting of an optional name and an action, which are performed using an Ansible module with specified arguments. Modules are reusable, standalone Python scripts that interact with the target host to execute specific tasks. For example, the ansible.builtin.shell module runs shell commands on the host. Ansible also supports blocks that group multiple tasks together.
Execution order. By default, Ansible executes tasks sequentially in the order in which they appear in the playbook. However, certain tasks can be configured to execute only when triggered by other tasks. These tasks, known as handlers, run in response to changes detected in preceding tasks.
Variables and Templates. Ansible uses variables to manage system differences, allowing the same variable name to hold different values for each host. These variables can be used in playbooks to adjust task behavior based on the target host. They can be defined in various ways, such as within playbooks or in Ansible’s inventory. Additionally, Ansible automatically gathers special variables called facts, which provide system information such as operating system details. Variable precedence is determined by the location where the variable is defined and follows a hierarchy of 22 rules (Ansible project contributors, 2025). Ansible leverages Jinja2 templating to enable dynamic expressions and access to variables and facts (Ansible project contributors, 2025), both within the playbooks and in configuration files used via the template module.
Push-based model. Unlike configuration management tools such as Chef and Puppet, Ansible follows a push-based configuration model. Typically, configuration management tools use a pull-based model that involves a server that stores configurations and target hosts that retrieve them. However, in a push model, the server actively sends configurations to the target hosts rather than waiting for them to request updates. Since target hosts do not initiate communication, push-based systems such as Ansible do not require an agent on each machine. Instead, a server, commanded by the Ansible user and with SSH access to all hosts, deploys Ansible modules to bring each host to the desired state.
Example. Figure 1 presents excerpts of an Ansible playbook for configuring a web server. This playbook consists of a single play. Here, we show three tasks and one handler. In line LABEL:line:hosts, the playbook specifies that it applies to all target hosts listed in the inventory. Line LABEL:line:roles defines the use of the webserver role, ensuring that all tasks within this role are executed. Line LABEL:line:fact demonstrates the use of an Ansible fact to display the operating system of the target host, using Jinja2 syntax for variable access, if the condition specified in line LABEL:line:condition is met. In line LABEL:line:notify, the playbook specifies that if the state of the nginx package changes, the handler Restart Nginx (line LABEL:line:restartnginx) should be triggered. Line LABEL:line:template employs a Jinja template to dynamically generate a default web page, allowing the use of Ansible variables for customization.
@@
[…]
7tasks:
8- name: Display system facts
9debug:
@@
13apt:
14name: nginx
[…]
18src: templates/index.html.j2
19dest: /var/www/html/index.html
20notify: Reload Nginx
[…]
22handlers:
[…]
24service:
25name: nginx
26state: restarted
27(...)
Figure 1. Excerpts from an Ansible playbook to configure a web server.
@@ 3. Mixed Methods Study
In this section, we outline our research questions and describe the methodology used to answer them. Our study follows a mixed methods approach (see Figure 2). First, we collected posts from Q&A platforms and used topic modeling to understand the challenges faced by Ansible practitioners. We then conducted semi-structured interviews to gain deeper insights into the participants’ experiences.
@@ 3.1. Research Questions
  • RQ1: [Issues] What are the common issues practitioners face when using Ansible? In this question, we investigate the common challenges that practitioners face while working with Ansible. By analyzing the emergent topics from practioners’ posts in Q&A platforms and practitioners feedback from interviews, we aim to categorize these challenges and identify patterns that can help us to understand pain points across various praticioner groups.
@@ 3.2. Topic Modeling
Building on previous studies (Begoug et al., 2023; Tanzil et al., 2023), we apply topic modeling techniques to identify the challenges practitioners face when using a specific tool, in our case, Ansible. In particular, we use TopicGPT, a prompt-based framework that uses large language models (LLMs) to uncover latent topics in a text collection (Pham et al., 2024). TopicGPT provides interpretable topics that align better with human categorizations than other approaches such as LDA and BERTopic (Pham et al., 2024). After data is collected, the TopicGPT framework involves three steps: Topic Generation, Topic Refinement, and Topic Assignment. In our study, we also perform manual validation of the Topic Assignment step. We configured TopicGPT to use GPT-4o mini via the OpenAI API, specifically the gpt-4o-mini-2024-07-18 version.
3.2.1. Data Collection
@@
We start by collecting posts about challenges, obstacles, and issues faced by Ansible practitioners from three Q&A platforms: StackOverflow (Stack Exchange, 2025b), Reddit (Reddit, 2025), and the Ansible Forum (Red Hat, 2025b) ( 1 in Figure 2). To collect posts from StackOverflow, we used the Stack Exchange API (Stack Exchange, 2025a) and retrieved all posts with the tag Ansible, which corresponds to a total of 23,000 posts. For Reddit, we used the PullPush API (PullPush team, 2025) to collect all 7,187 posts on the subreddit r/ansible that contained the keyword Ansible in addition to keywords possibly related to issues, such as error, help, or slow. For the Ansible Forum, we used web scraping techniques to collect 28,970 posts in the Get Help and Archives categories. We included the category Archives in our analysis, as it contains all posts from the Ansible Google Group, the predecessor of the Ansible Forum, which were later migrated to the Ansible Forum (Red Hat, 2025c). We filtered out posts related to announcements and releases by the Ansible team. In total, we collected 59,157 posts from the three Q&A platforms, forming our final dataset. Data collection took place between January 24, 2025, and February 1, 2025.
3.2.2. Topic Generation
@@
We use TopicGPT to generate a set of topics given our whole dataset as input ( 2 in Figure 2). A topic consists of a name and a description. For each post, TopicGPT prompts GPT-4o Mini to assign one or more of the already generated topics and generate any new relevant topics that are missing, assigning them accordingly. To increase the probability that the topics generated are relevant to our study, we added to the default prompt of TopicGPT the phrase: “The topics must be related to challenges, issues, or problems that developers face while using Ansible.”. In the end, TopicGPT generated 3,573 topics ( 3 in Figure 2).
3.2.3. Topic Refinement
TopicGPT produces numerous topics that, while generally relevant, are often overly granular, duplicated, or out of scope (Pham et al., 2024). To mitigate these limitations, it incorporates an automated topic refinement step. However, in our first experiments, we achieved unsatisfactory results, even when using more powerful LLMs such as o3-mini-2025-01-31. For example, in one of our test runs, this approach was unable to merge the following pair of topics with a cosine similarity of $0.9600$, which TopicGPT uses to measure semantic proximity:
1Regex Issues Desc: Problems related to the use of regular expressions in Ansible tasks.
2Regex Desc: Issues related to the use of regular expressions in Ansible.
@@ 3.2.3. Topic Refinement
For this reason, we adopted an approach similar to TopicGPT, but instead of using LLMs, we relied on manual effort to ensure a more reliable final topic list ( 4 in Figure 2). First, instead of merging the topics first and then removing the infrequent ones, we reversed the order. This approach reduced the number of topic pairs that needed to be analyzed during the merging process. The topic distribution follows a power law, where a small number of topics account for the majority of post attributions, whereas most topics are assigned to only a few posts. To ensure adequate coverage, we initialized an empty set and incrementally added topics in descending order of attribution frequency until reaching 80% of the total attributions. Topics that were not included in the final set were discarded. After this step, our topic list was reduced from 3,573 to 153 topics.
@@
Next, we applied an iterative manual refinement process similar to the merging step used in TopicGPT. In each iteration, we systematically evaluated topic pairs, starting with those exhibiting the highest cosine similarity and proceeding to those with the lowest similarity down to a predefined threshold for that iteration. Similarly to TopicGPT, we used Sentence-Transformer embeddings (Reimers and Gurevych, 2019) to calculate cosine similarity. Then, for each pair, one author determined whether the topics should be merged and, if so, whether they should be combined into one of the existing topics or assigned to a newly created topic. The criterion for merging topics was the presence of significant semantic overlap between them. If a topic was merged in a given iteration, pairs that included the merged topic were excluded from further evaluation within that iteration. A second author independently reviewed the merged topics, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. The process was repeated six times until no further merges were identified, resulting in a candidate topic list. In the first three iterations, we applied a cosine similarity threshold of 0.7, while in the last three iterations, we lowered the threshold to 0.6. After the merging process, we reduced our topic list from 153 to 105 topics.
@@
Finally, following the approaches used in previous studies (Asmussen and Møller, 2019; Jacobi et al., 2018), we manually inspected the topics to identify those relevant to our study. A topic was considered relevant if it is related to the challenges, frustrations, issues, or drawbacks associated with using Ansible. We considered only topics that were specific enough to provide meaningful information about a post. For example, this led us to remove the topic Task Execution with description “Issues related to the execution of tasks in Ansible, including conditions and dependencies for task execution.”. In addition, we refined topic names and descriptions. A topic name was considered improved if it more effectively represented its description, while a description was considered improved if it maintained its core ideas while increasing generality, eliminating unnecessary details, or using a clearer and more concise language. Finally, if we identified any additional pairs of topics that met our merging criteria but were not initially detected during the merging process, we merged them accordingly.
@@
An author reviewed each topic, determining whether it should be removed, renamed, or modified. Subsequently, a second author independently reviewed the proposed changes, documenting any disagreements or ambiguous cases requiring further discussion. After this independent review, both authors met to resolve discrepancies and reach a consensus. This resulted in a final list of 87 topics ( 5 in Figure 2, see Supplemental Materials for the full topic list).
3.2.4. Topic Assignment
We use the topic list generated in Section 3.2.3 and TopicGPT to assign topics to each post. We also apply the self-correction step provided in the TopicGPT framework (Pham et al., 2024) ( 6 in Figure 2).
3.2.5. Validation
@@
To assess the quality of the topic assignment described in Section 3.2.4, we employ the confirmation labeling approach (Yang et al., 2014). To calculate the precision of topic assignment, this method involves sampling (post, topic) pairs of assignments under evaluation. Rather than presenting raters with the entire taxonomy and asking them to manually assign topics to posts, confirmation labeling simplifies the process by requiring raters to provide binary judgments on whether a given topic is relevant to a particular post. This approach effectively reduces the cognitive overload of a large and complex taxonomy (Yang et al., 2014).
Two authors independently performed confirmation labeling on a sample of 300 (post, topic) pairs ( 7 in Figure 2). The inter-rater agreement, measured using Cohen’s Kappa (Cohen, 1960), was approximately 0.29, which according to Landis and Koch’s interpretation (Landis and Koch, 1977), is classified as a fair agreement. A third author performed confirmation labeling on the (post, topic) pairs where the initial two authors’ judgments diverged. The final judgment was decided by majority vote. Ultimately, TopicGPT’s topic assignment achieved a precision ( 8 in Figure 2) of approximately 82.67%.
3.2.6. Analysis
@@
To draw relevant insights across our topic analysis and the interviews, we merged the final list of topics derived from TopicGPT with the thematic codes identified in the interview data (Section 3.3). Specifically, we integrated the topics with a set of Challenges codes by aligning each topic with the most closely related theme. This process was iterative and followed a double-blind procedure: first, two coders independently grouped the topics into themes. Then, both coders met to discuss disagreements and reach a consensus. In analyzing the merged data, we examined the relative frequency of each topic across the resulting themes and investigated how these frequencies evolved. We restricted our scope to the last five years of Q&A data to better capture current and emerging concerns, totaling 23,215 posts. In addition to recent posts, we leveraged the data older than five years to identify evolution trends.

### v3

@@ Abstract.
Infrastructure as Code (IaC) tools have transformed the way IT infrastructure is automated and managed, but their growing adoption has also exposed numerous challenges for practitioners. In this paper, we investigate these challenges through the lens of Ansible, a popular IaC tool. Using a mixed methods approach, we investigate challenges faced by practitioners. We analyze 59,157 posts from Stack Overflow, Reddit, and the Ansible Forum to identify common pain points, complemented by 20 semi-structured interviews with practitioners of varying expertise levels.
Based on our findings, we highlight key directions for improving Ansible, with implications for other IaC technologies, including stronger failure locality to support debugging, clearer separation of language and templating boundaries, targeted documentation, and improved execution backends to address performance issues. By grounding these insights in the real-world struggles of Ansible users, this study provides actionable guidance for tool designers and for the broader IaC community, and contributes to a deeper understanding of the trade-offs inherent in IaC tools.
Infrastructure as Code, Ansible, Practitioners Issues, User Study, Topic Modeling
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: International Conference on Evaluation and Assessment in Software Engineering; 9 – 12 June, 2026; Glasgow, United Kingdomisbn: 978-1-4503-XXXX-X/2018/06copyright: none
@@ 1. Introduction
In this paper, we aim to deepen the understanding of the challenges faced by Ansible users. Ansible is currently the most widely used configuration management tool (Stack Overflow, 2024) and, although Ansible is known to have issues, such as difficulties with variable scoping (Opdebeeck, 2024) and debugging (Hacker News User gizdan, 2021), the specifics of these challenges and their impact on practitioners remain underexplored. Using a mixed methods approach, we analyze a large corpus of online discussions and conduct semi-structured interviews to systematically examine the obstacles faced by Ansible users. Our study is the first to provide a detailed account of these issues and to offer actionable recommendations for improvement. Our contributions are:
@@
  • Recommendations for improving Ansible, with implications for other IaC technologies, focusing on stronger failure locality to support debugging, better defined language boundaries, targeted documentation, and improved execution backends.
Our study aims to inform Ansible’s developers, educators, and the broader IaC community, fostering improvements that enhance the effectiveness and usability of IaC tools. A replication package with all the supplementary materials, data, and code used in this study is available (Anonymous Author(2025), s).
@@ 2. Ansible
Ansible is the most widely used configuration management tool (Stack Overflow, 2024). Configurations are written as YAML playbooks composed of plays and tasks, optionally organized into reusable roles. Tasks call modules with arguments to enforce the desired state on target hosts. For example, the ansible.builtin.shell module runs shell commands on the host. Ansible executes tasks sequentially in the order in which they appear in a playbook. It uses variables to manage system differences, allowing the same variable name to hold different values for each host. These variables can be used in playbooks to adjust task behavior based on the target host. Ansible leverages Jinja2 templating to enable dynamic expressions and access to variables (Ansible project contributors, 2025), both within the playbooks and in configuration files used via the template module. Unlike pull-based technologies such as Chef or Puppet, Ansible follows an agentless, push-based model, typically applying changes from a control node to target hosts over SSH.
Fig. 1 shows excerpts of an Ansible playbook for the configuration of a web server, consisting of a single play with three tasks. The playbook targets all hosts in the inventory, uses Ansible variables and Jinja2 expressions for conditional execution, and applies a template to generate a customized web page.
@@
6tasks:
7- name: Display system facts
8debug:
@@
11- name: Ensure Nginx is installed
12apt:
13name: nginx
14- name: Deploy an index.html template
@@
16src: templates/index.html.j2
17dest: /var/www/html/index.html
18(...)
Figure 1. Excerpt from a web server Ansible playbook.
@@ 3. Mixed Methods Study
In this section, we outline our research questions and describe the methodology used to answer them. Our study follows a mixed methods approach (see Fig. 2). First, we collected posts from Q&A platforms and used topic modeling to understand the challenges faced by Ansible practitioners. We then conducted semi-structured interviews to gain deeper insights into the participants’ experiences.
@@ 3.1. Research Questions
  • RQ1: [Issues] What are the common issues practitioners face when using Ansible? In this question, we investigate the common challenges that practitioners face while working with Ansible. By analyzing the emergent topics from practitioners’ posts in Q&A platforms and practitioners feedback from interviews, we aim to categorize these challenges and identify patterns that can help us to understand pain points across various praticioner groups.
@@
RQ1 is addressed using both topics extracted from online Q&A data and practitioner interviews, enabling triangulation between large-scale community discussions and in-depth qualitative insights. RQ2 and RQ3 are answered exclusively through interviews, as public technical artifacts primarily reflect post-adoption troubleshooting and rarely capture adoption rationales or broader improvement considerations. The key directions discussed in Section 7 are grounded in the combined results of all three research questions, integrating evidence from both data sources.
@@ 3.2. Topic Modeling
Building on previous studies (Begoug et al., 2023; Tanzil et al., 2023), we apply topic modeling techniques to identify the challenges practitioners face when using a specific tool, in our case, Ansible. We use TopicGPT, a prompt-based framework that uses LLMs to uncover latent topics in a text collection (Pham et al., 2024). TopicGPT provides interpretable topics that align better with human categorizations than other topic modeling approaches (Pham et al., 2024). After data is collected, TopicGPT consists of three steps: Topic Generation, Topic Refinement, and Topic Assignment. We also perform manual validation of the Topic Assignment step. We configured TopicGPT to use GPT-4o mini via the OpenAI API.
Data Collection
@@
We start by collecting posts about challenges, obstacles, and issues faced by Ansible practitioners from three Q&A platforms: StackOverflow (Stack Exchange, 2025b), Reddit (Reddit, 2025), and the Ansible Forum (Red Hat, 2025a) ( 1 in Fig. 2). To collect posts from StackOverflow, we used the Stack Exchange API (Stack Exchange, 2025a) and retrieved all posts with the tag Ansible, which corresponds to 23,000 posts. For Reddit, we used the PullPush API (PullPush team, 2025) to collect all 7,187 posts on the subreddit r/ansible that contained the keyword Ansible in addition to keywords possibly related to issues, such as error and help. For the Ansible Forum, we used web scraping techniques to collect 28,970 posts in the Get Help and Archives categories. We included the category Archives in our analysis, as it contains all posts from the Ansible Google Group, the predecessor of the Ansible Forum, which were migrated to the Ansible Forum (Red Hat, 2025b). We filter out posts related to announcements by the Ansible team. In total, we collected 59,157 posts from the three Q&A platforms, forming our final dataset. Data collection took place between January 24 and February 1, 2025.
Topic Generation
@@
We use TopicGPT to generate a set of topics given our whole dataset as input ( 2 in Fig. 2). A topic consists of a name and a description. For each post, TopicGPT prompts GPT-4o Mini to assign one or more of the already generated topics and generate any new relevant topics that are missing, assigning them accordingly. To ensure relevance, we constrained the TopicGPT prompt to generate topics centered on the challenges faced by Ansible practitioners. In the end, TopicGPT generated 3,573 topics ( 3 in Fig. 2).
Topic Refinement
@@
TopicGPT produces numerous topics that, while generally relevant, are often overly granular, duplicated, or out of scope (Pham et al., 2024). To mitigate these limitations, we include an automated topic refinement step. However, in our preliminary experiments, automated refinement did not reliably merge semantically equivalent topics. For this reason, we adopted an approach similar to TopicGPT, but instead of using LLMs, we relied on manual effort to ensure a reliable final topic list ( 4 in Fig. 2). First, instead of merging the topics first and then removing the infrequent ones, we reversed the order, reducing the number of topic pairs that needed to be analyzed during the merging process. The topic frequencies follow a power-law distribution. To ensure coverage, we retained the most frequent topics until they accounted for 80% of all attributions. Topics that were not included in the final set were discarded. After this step, our topic list was reduced from 3,573 to 153 topics.
@@
Next, we applied an iterative manual refinement process similar to the merging step in TopicGPT. Each iteration evaluated topic pairs above a predefined cosine similarity threshold. Then, for each pair, one author determined whether the topics should be merged and, if so, whether they should be combined into one of the existing topics or assigned to a newly created topic. The criterion for merging topics was the presence of significant semantic overlap between them. A second author reviewed the merged topics, documenting any disagreements or ambiguous cases requiring further discussion. Subsequently, both authors met to resolve discrepancies and reach a consensus. The process was repeated six times until no further merges were identified, resulting in a candidate topic list. After the merging process, we reduced our topic list from 153 to 105 topics.
@@
Finally, following the approaches used in previous studies (Asmussen and Møller, 2019; Jacobi et al., 2018), we manually inspected the topics to identify those relevant to our study. A topic was considered relevant if it is related to the challenges, frustrations, issues, or drawbacks associated with using Ansible. We considered only topics that were specific enough to provide meaningful information about a post. In addition, we refined topic names and descriptions. A topic name was considered improved if it more effectively represented its description, while a description was considered improved if it maintained its core ideas while increasing generality, eliminating unnecessary details, or using a clearer and more concise language. Finally, we merged any additional pairs of topics that met our merging criteria but were not initially detected during the merging process.
@@
An author reviewed each topic, determining whether it should be removed, renamed, or modified. Subsequently, a second author independently reviewed the proposed changes, documenting any disagreements or ambiguous cases requiring further discussion. Subsequently, both authors met to resolve discrepancies and reach a consensus. This resulted in a final list of 87 topics ( 5 in Fig. 2, see Supplemental Materials for the full topic list).
Topic Assignment
We used TopicGPT to assign the final topics to posts and applying its self-correction mechanism (Pham et al., 2024) ( 6 in Fig. 2).
Validation
@@
We evaluated topic assignment using confirmation labeling (Yang et al., 2014). To calculate the precision of topic assignment, this method involves sampling (post, topic) pairs of assignments under evaluation. Rather than presenting raters with the entire taxonomy and asking them to manually assign topics to posts, confirmation labeling simplifies the process by requiring raters to provide binary judgments on whether a given topic is relevant to a particular post. This approach effectively reduces the cognitive overload of a large and complex taxonomy (Yang et al., 2014). Two authors independently performed confirmation labeling on a sample of 300 (post, topic) pairs ( 7 in Fig. 2). A third author resolved disagreements between the initial raters, with final decisions made by majority vote, yielding a precision of 82.67% ( 8 in Fig. 2).
Analysis
@@
To integrate topic modeling and interview findings, we merged the final topic list with the thematic codes identified in the interviews (Section 3.3). Specifically, we integrated the topics with a set of Challenges codes by aligning each topic with the most closely related theme. This process was iterative and followed a double-blind procedure: first, two coders independently grouped the topics into themes. Then, both coders met to discuss disagreements and reach a consensus. In analyzing the merged data, we examined the relative frequency of each topic across the resulting themes and investigated how these frequencies evolved. To better reflect current and emerging concerns, we limited our analysis to Q&A data from the five years preceding the end of data collection on February 1, 2025, resulting in a total of 24,132 posts. In addition to recent posts, we leveraged the data older than five years to identify evolution trends.